throught (-1, -4), perpendicualr to 7x-3y=5
Answer by scott8148(6628)

swap the x and y coefficients and change one of the signs to get the perpendicular

plug in the given point to find the new constant term

3x + 7y = ? ___ 3(-1) + 7(-4) = -31

3x + 7y = -31